OpenStudy (anonymous):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4-x and y = 2x + 3 intersect are the solutions of the equation 4-x = 2x + 3. (4 points) Part B: Make tables to find the solution to 4-x = 2x + 3. Take the integer values of x between -3 and 3. (4 points) Part C: How can you solve the equation 4-x = 2x + 3 graphically? (2 points)

OpenStudy (anonymous):

I have everything else besides Part B.. Could you please help me with part b?

OpenStudy (anonymous):

@dan815 @ganeshie8 @StudyGurl14

OpenStudy (stamp):

b. 4 -x = 2x + 3 1 = 3x x = 1/3

OpenStudy (stamp):

Table for b.

OpenStudy (anonymous):

Thats all? isnt that just solving the equation? like what is the integer values of x between -3 and 3.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@hartnn Please help me :'(

OpenStudy (anonymous):

@pizza12 can you help me please?

OpenStudy (anonymous):

sorry i forgot how to do this :'(

OpenStudy (anonymous):

Actually nvm

OpenStudy (anonymous):

Stamp posted a good answer i just didnt see it at first.